# CushionCatch: Compliant Catching Mechanism for Mobile Manipulators via Combined Optimization and Learning

### v1

@@ IV-C1 PRC Planner
In the high-level capture planner, it provide the joint configuration of the robot to catch the ball. For the catching task, ensuring both speed and smoothness in trajectory planning is essential. Considering the maximum acceleration and speed limits for each joint, we determine the minimum trajectory planning time for each joint, denoted as $t_{prc}$:
@@
$t_{prc}=\begin{cases}2\sqrt{\frac{\triangle q}{\ddot{q}_{max}}},&\quad\triangle q<2\triangle q_{acc}\\ 2\frac{\dot{q}_{max}}{\ddot{q}_{max}}+\frac{\triangle q-\ddot{q}_{max}\left(\frac{\dot{q}_{max}}{\ddot{q}_{max}}\right)^{2}}{\dot{q}_{max}},&\quad\triangle q\geq 2\triangle q_{acc}\end{cases}$ (11)
@@
where $\triangle q=|q_{ca}-q_{0}|$, $q_{acc}=\frac{1}{2}\ddot{q}_{max}\left(\frac{\dot{q}_{max}}{\ddot{q}_{max}}\right)^{2}$. This calculation assumes that the joint accelerates to maximum speed using the maximum acceleration and then maintains that speed until the deceleration phase. While such an extreme motion is not used in practice, this simplified method helps us quickly estimate the duration of the joint movement. Then, we can define the time parameter $T_{prc}$ of PRC as follow:
$T_{prc}=\lambda\cdot max(t_{prc}^{i}),\quad i\in[1,n]$ (12)
@@
where $\lambda>1$ is a proportion parameter to enlarge planning time appropriately. Then, we can get the PRC planner in (6):
${\mathit{\Phi}}_{prc}=\text{Poly5}(t;T_{prc},\boldsymbol{q}_{0},\boldsymbol{q}_{ca}))$ (13)
@@
where Poly5() represents a Quintic Polynomial Planning that maps the time $t$(ranging from 0 to $T_{prc}$) to the joint position based on the inputs $T_{prc},\boldsymbol{q}_{0},\boldsymbol{q}_{ca}$.
IV-C2 POC Planner
After the robot catches the ball, it needs to apply compliant control to have a cushioning mechanism. Given the velocity-compliant sequence $\boldsymbol{\psi}$ produced by the network, we need to track this sequence. Additionally, to ensure safety, we must consider potential collisions between the mobile base and the manipulator, as well as between the manipulator and the ground. Therefore, we look at how Control Barrier functions(CBF) can encode safety constraints. The POC planner can be expressed as follows:
${\mathit{\Phi}}_{poc}=\min_{\dot{\boldsymbol{q}}}\frac{1}{2}\left(\|\dot{\boldsymbol{q}}\|^{2}+\|\boldsymbol{\delta}\|^{2}\right)$ (14a)
[…]
$\dot{f}(z_{c})\cdot{}_{c}^{w}\boldsymbol{J}_{z}\cdot\dot{\boldsymbol{q}}\geq-\gamma_{z}f(z_{c})$ (14c)
$\dot{g}(x_{c},y_{c})\cdot{}_{c}^{w}\boldsymbol{J}_{xy}\cdot\dot{\boldsymbol{q}}\geq-\gamma_{xy}g(x_{c},y_{c})$ (14d)
where $\psi_{i}$ represents the $i$-th sequence of $\boldsymbol{\psi}$ and $\boldsymbol{\delta}$ is the slack vector. The components of the Jacobian in $z$-axis and the $x,y$-axis are denoted by ${}_{c}^{w}\boldsymbol{J}_{z}$ and ${}_{c}^{w}\boldsymbol{J}_{xy}$. The function $f(z_{c})$ represents the safe distances between the container and the ground. And $g(x_{c},y_{c})$ is the safe horizontal distance between the container and the mobile base. Safety is maintained when $f(z_{c})>0$ and $g(x_{c},y_{c})>0$. In the optimization problem above, equality constraints ensure the container keeps track of expectations $\boldsymbol{\psi}$. To ensure safety, Control Barrier Functions (CBFs) are used as inequality constraints, which help maintain safe forward invariance [28, 29]. Although the compliant sequence learned from human demonstrations may indicate potential collisions, these are effectively mitigated through the implementation of safety constraints. By imposing limits on the container’s height along the z-axis and regulating its horizontal distance from the base, the optimization process filters out potential collisions while minimizing deviations from the desired trajectory. This ensures safety without compromising performance. The collision scenarios are shown in Figure 3.
@@ IV-D Compliant Sequence Learning
[…]
@@
$input_{i}=input_{i-1}\oplus out_{i-1}+PE(i-1,6)$ (15a)
$out_{i}=\boldsymbol{G}(input_{i})$ (15b)
where $PE(\cdot)$ means positional encoding and $i$ is the $i$-th time step, $\oplus$ represents concatenation vertically and $\boldsymbol{G}(\cdot)$ is the one-step-LSTM module.
@@
$pe(l,2i,k)=\sin\left(\frac{l}{10000^{2i/k}}\right)$ (16a)
$pe(l,2i+1,k)=\cos\left(\frac{l}{10000^{2i/k}}\right)$ (16b)
$PE(l,k)={pe(l,0,k)|pe(l,1,k)|...|pe(l,k,k)}$ (16c)
@@
Figure 6: The robot motions under CCCM framework to catch a ball with compliant control. They represent the starting state of motion, the end state of PRC planner, and the end state of POC planner.
@@ V EXPERIMENTS
We evaluate our framework through experiments conducted both in simulation and on a physical mobile manipulator. We use the following value for the parameters of our framework: $k_{ad}=0.0295$ in (3); $\alpha=2$ in (9a); $\beta=0.5$ in (9d); ${\mathit{\Lambda}}_{b}=5,{\mathit{\Lambda}}_{m}=1$ in (10); $\lambda=1.5$ in (12); $\gamma_{z}=0.1$ in (14c); $\gamma_{xy}=0.1$ in (14d). In P-LSTM, we set the hidden size to 64 and a fully connected layer block from 64 to 6. During training, we use mean square error(MSE) between predicted output and ground truth value, the learning rate is 0.001 and the max output sequence length is capped at 16.
@@ V-A Data Collection and Processing
We capture the velocity and position sequences of the ball using a NOKOV motion-capture device during human-catching demonstrations. These sequences cover a short period from just before the ball is caught until it comes to a complete stop. We identify the moment of catching by detecting the point with the highest average velocity variation across the $x,y,z$-axes. This point is then used to split the sequence into input data and corresponding output labels.
V-B Simulation Scenarios
We use the Swift and Rtb Python libraries [31] as our simulated environments. Initially, we generated 6,000 ball samples $[p_{x},p_{y},p_{z},\dot{p}_{x},\dot{p}_{y},\dot{p}_{z}]$, restricting their positions within a cylindrical space of a specified height. We ensured that their trajectories, for $t>0.5$, would pass through the robot’s capture range. Then ball parameters are generated randomly by procedure. Additionally, we performed ablation studies on the collision avoidance constraints in (14c) and (14d) during the POC phase (2,000 pieces of data for each test). The statistical results are presented in Table I.
TABLE I: Simulation Results Over 6000 Random Catching Tasks
@@ V-B Simulation Scenarios
Fig. 5 shows the robot motions of some experiments where the ball comes from different directions. Table I demonstrates the results of 6000 random catching tasks, showing that the CCCM framework achieves a 98.70% success rate for catching. In practice, constraints (14c) and (14d) do not affect the PRC planner, resulting in similar ”Not Catch” rates. However, incorporating collision avoidance constraints notably reduces the incidence of collisions. Without (14c) or (14d), the probability of collisions in their respective dimensions increases substantially.
Figure 7: The change of the joint angles of the robot during a catch, the expected joint angles calculated by the high-level planner, and the trajectory of the ball and the container. The catching time is $t_{ca}=0.598$.
Figure 8: Manipulator joint torques without and with the POC planner. The POC planner reduces the impact torques when robot catching.
V-C Real-World Scenarios
@@
As shown in Fig. 6, a real-world experiment is conducted on a mobile manipulator. We use a C-100 differential mobile base($n_{b}$ = 2) and a franka panda manipulator($n_{m}$ = 7), interfacing with the robot through ROS Noetic. NOKOV motion capture provides position and velocity measurements, which are used for both ball state estimation and mobile base state tracking. We use a standard tennis ball fitted with a reflective strip for tracking purposes. When a human hand throws the ball, a Kalman filter models the ball’s dynamics using feedback from the initial frames and updates the model in the subsequent few frames. This model is then used to predict the ball’s trajectory over the next 1.5s. Notably, due to the coupling effects of differential wheel steering, our high-level planner considers only the combined motion $q_{1}$ and restricts the differential motion $q_{0}$ to 0 to achieve more accurate tracking in the PRC. We conducted a total of 100 hand-throwing experiments. However, due to inaccuracies in hand throwing, 19 attempts clearly fell outside the robot’s capture range. Out of the remaining 81 valid throws, the robot successfully captured the ball 75 times, resulting in a success rate of 92.59%, without any collision. The failure occurred because noise in the capture system caused a discrepancy between the filter’s estimated state and the actual state, leading to deviations in the predicted trajectory. In related systems [1, 20], large-scale data simulations were not conducted, yet we still outperformed them in real-world experiments. Additionally, we incorporated compliant behavior, making the system safer and more relevant.
Fig. 7(a) illustrates the expected output of the high-level capture planner and the changes in joint angles during the catching. The PRC planner successfully guided all joints to meet the expected positions before the capture time $t_{ca}=0.598$ and the system transitioned smoothly into the POC planner phase. Fig. 7(b) shows the actual, predicted trajectory of the ball and the trajectory of the container during catching. The trajectory prediction is accurate. It is worth noting that the position of the container does not coincide with the ball’s position after the catch due to the depth of the net bag. Fig. 8(a) and Fig. 8(b) present the joint torque data without and with the POC planner, respectively. $q_{3}$ and $q_{5}$ are the main joints that receive impact in the process of catching. Clearly, the addition of the POC planner to have compliant catching significantly reduced the impact on the joints. By comparing the maximum torque of $q_{3},q_{5}$, the compliant strategy reduces the impact torques by an average of 33.2%, demonstrating that the compliant strategy learned by the network is effective. At the same time, normal torque fluctuations in other joints are expected due to joint movements during the POC planner. Moreover, it is worth mentioning that there is no collision in the physical experiments. The safety is ensured through our POC planner when applying the compliant sequences learned from human demonstrations to the robot.

### v3

@@ IV-C PRC Planner
In the high-level capture planner, the robot’s joint configuration $\bm{q}_{ca}$ for catching the ball is determined. The robot then needs to reach this joint configuration as quickly as possible. Considering the maximum acceleration and speed limits for each joint, we determine the minimum trajectory planning time for each joint, denoted as $t_{prc}$:
@@
$t_{prc}=\begin{cases}2\sqrt{\frac{\triangle q}{\ddot{q}_{max}}},&\quad\triangle q<2q_{acc}\\ 2\frac{\dot{q}_{max}}{\ddot{q}_{max}}+\frac{\triangle q-\ddot{q}_{max}\left(\frac{\dot{q}_{max}}{\ddot{q}_{max}}\right)^{2}}{\dot{q}_{max}},&\quad\triangle q\geq 2q_{acc}\end{cases}$ (7)
@@
where $\triangle q=|q_{ca}-q_{0}|$, $q_{acc}=\frac{1}{2}\ddot{q}_{max}\left(\frac{\dot{q}_{max}}{\ddot{q}_{max}}\right)^{2}$. This section determines the shortest theoretical time for the robot to reach the target configuration $q_{ca}$ if following the limits on maximum acceleration and speed. While such an extreme motion is not used in practice, this simplified method helps us quickly estimate the duration of the joint movement. we can define the time parameter $T_{prc}$ of PRC as follow:
$T_{prc}=\lambda\cdot max(t_{prc}^{i})\quad i\in[1,n],$ (8)
@@
where $\lambda>1$ is a proportion parameter to enlarge planning time appropriately. Then, we can get the PRC planner $\phi_{prc}$ in joint space:
${\mathit{\Phi}}_{prc}=\text{Poly5}(t;T_{prc},\bm{q}_{0},\bm{q}_{ca})),$ (9)
@@
where Poly5() represents a Quintic Polynomial Planning that maps the time $t$(ranging from 0 to $T_{prc}$) to the joint position based on the inputs $T_{prc},\bm{q}_{0},\bm{q}_{ca}$.
@@ IV-D Compliant Sequence Learning
Given the velocity and position states of the object before catching, we employ the P-LSTM network to output the compliant sequences required for POC phase.
@@
$input_{i}=input_{i-1}\oplus out_{i-1}+PE(i-1,6),$ (10a)
$out_{i}=\mathcal{L}(input_{i}),$ (10b)
where $PE(\cdot)$ means positional encoding and $i$ is the $i$-th time step, $\oplus$ represents concatenation vertically and $\mathcal{L}(\cdot)$ is the one-step-LSTM module.
@@
$pe(l,2i,k)=\sin\left(\frac{l}{10000^{2i/k}}\right),$ (11a)
$pe(l,2i+1,k)=\cos\left(\frac{l}{10000^{2i/k}}\right),$ (11b)
$PE(l,k)={pe(l,0,k)|pe(l,1,k)|...|pe(l,k,k)},$ (11c)
@@
The P-LSTM network is trained on data obtained from human demonstrations. During the application phase, the predicted ball position and velocity at $\bm{q}_{ca}$ , denoted as $\hat{{\bm{p}}}(\bm{q}_{ca})$ and $\hat{\dot{\bm{p}}}(\bm{q}_{ca})$, is computed and subsequently used as input to the network. Based on this input, the network generates the compliant velocity sequence $\bm{\psi}\in\mathbb{R}^{6\times s}$ , where $s$ denotes the length of the sequence. We focus on tracking the compliant velocity sequence because position is heavily dependent on the world coordinate system. In contrast, velocity is relative and more general, independent of the position of world coordinate system, making it more universal.
IV-E POC Planner
After the robot catches the ball, it needs to apply compliant control to have a cushioning mechanism. Given the velocity-compliant sequence $\bm{\psi}$ produced by the network, we need to track this sequence. Additionally, applying the sequence directly to the robot could result in a series of collisions due to the structural differences between the robot and the human, as shown in Fig. (4). To ensure safety, we must account for potential collisions between the mobile base and the manipulator, as well as between the manipulator and the ground. Therefore, we look at how Control Barrier functions(CBF) [31] can encode safety constraints. The POC planner can be expressed as follows:
${\mathit{\Phi}}_{poc}=\min_{\dot{\bm{q}}}\frac{1}{2}\left(\|\dot{\bm{q}}\|^{2}+\|\bm{\delta}\|^{2}\right)$ (12a)
[…]
$\dot{f}(z_{c})\cdot{}_{c}^{w}\bm{J}_{z}\cdot\dot{\bm{q}}\geq-\gamma_{z}f(z_{c}),$ (12c)
$\dot{g}(x_{c},y_{c})\cdot{}_{c}^{w}\bm{J}_{xy}\cdot\dot{\bm{q}}\geq-\gamma_{xy}g(x_{c},y_{c}),$ (12d)
where $\psi_{i}$ represents the $i$-th sequence of $\bm{\psi}$ and $\bm{\delta}$ is the slack vector to reduce the rigor of the equation. The components of the Jacobian in $z$-axis and the $x,y$-axis are denoted by ${}_{c}^{w}\bm{J}_{z}$ and ${}_{c}^{w}\bm{J}_{xy}$. The function $f(z_{c})$ represents the safe distances between the container and the ground. And $g(x_{c},y_{c})$ is the safe horizontal distance between the container and the mobile base. Safety is maintained when $f(z_{c})>0$ and $g(x_{c},y_{c})>0$. In the optimization problem above, equality constraints ensure the container keeps track of expectations $\bm{\psi}$. To ensure safety, CBFs are used as inequality constraints, which help maintain safe forward invariance [32, 33], that is $f(z_{c})>0$ and $g(x_{c},y_{c})>0$. By imposing limits on the container’s height along the z-axis and regulating its horizontal distance from the base, the optimization process filters out potential collisions while minimizing deviations from the desired trajectory.
Figure 5: Collecting compliant data from human demonstration in the real world
Figure 6: Training loss curves for compliant sequence learning across different networks.
@@ IV-E POC Planner
Figure 8: The robot motions under CCMM framework to catch a ball with compliant control. They represent the starting state of motion, the end state of PRC planner, and the end state of POC planner.
@@ V EXPERIMENTS
We evaluate our framework through experiments conducted both in simulation and on a physical mobile manipulator. We use the following value for the parameters of our framework: $k_{ad}=0.0295$ in (3); $\alpha=2$ in (5a); $\beta=0.5$ in (5d); ${\mathit{\Lambda}}_{b}=5,{\mathit{\Lambda}}_{m}=1$ in (6); $\lambda=1.5$ in (8); $\gamma_{z}=0.1$ in (12c); $\gamma_{xy}=0.1$ in (12d). We use ”SLSQP” method to solve nonlinear optimization problems (5a) and (12a). In P-LSTM, we set the hidden size to 64 and a fully connected layer block from 64 to 6. During training, we use mean squared error (MSE) between the predicted output and the ground truth. The learning rate is set to 0.001, and the max output sequence length is capped at 16.
@@ V-A Data Collection and Processing
We capture the velocity and position sequences of the ball using a NOKOV motion-capture device during human-catching demonstrations as Fig. (5) shows. Reflective balls are placed on both the glove and the ball. We use the distance between them to determine when to start recording data. In total, 302 trajectories were collected. To distinguish between inputs and outputs of the data for the network, we identified the point with the highest average velocity variation across all data points. This point is then used to split the sequence into input data and the corresponding output labels.
V-B Network Comparison Experiment
To evaluate our P-LSTM method, we compare it with four other methods: LSTM without PE, transformer without PE, transformer with PE, and diffusion policy. All methods use 2 hidden layers with a hidden size of 32, and the transformer’s number of heads is set to 2 (the input size is 6, with 3 position components and the remaining 3 being velocities, so 2 heads are sufficient to capture the relationships between them). The training and testing metrics are the MSE between the six-dimensional true sequences and the predicted sequences. We randomly divide the data into a training set (4/5) and a test set (1/5).
As shown in TABLE I and Fig. 6, transformer-based methods struggle to converge and validate effectively due to their high parameter demands and the short input sequences. The diffusion policy also fails to train effectively due to the limited data size. Additionally, adding PE to both LSTM and transformer models reduces the error, with almost no impact on inference speed in test experiments.”
V-C Simulation Scenarios
We use the Swift and Rtb Python libraries [34] as our simulated environments. Initially, we generated 6,000 ball samples $[p_{x},p_{y},p_{z},\dot{p}_{x},\dot{p}_{y},\dot{p}_{z}]$, restricting their positions within a cylindrical space of a specified height. We ensured that their trajectories, for $t>0.5$, would pass through the robot’s capture range. Then ball parameters are generated randomly by procedure. Additionally, we performed ablation studies on the collision avoidance constraints in (12c) and (12d) during the POC phase (2,000 pieces of data for each test). The statistical results are presented in TABLE II.
TABLE I: Test Experiments
Method Test MSE Inference time (s)
Transformer 0.168 0.201
Transformer+PE 0.152 0.223
LSTM 0.096 0.061
LSTM+PE 0.091 0.062
Diffusion 0.161 0.307
TABLE II: Simulation results over 6000 random catching tasks
@@ V-C Simulation Scenarios
Fig. 7 shows the robot motions of some experiments where the ball comes from different directions. TABLE II demonstrates the results of 6000 random catching tasks, showing that the CCMM framework achieves a 98.70% success rate for catching. In practice, constraints (12c) and (12d) do not affect the PRC planner, resulting in similar ”Not Catch” rates. However, incorporating collision avoidance constraints notably reduces the incidence of collisions. Without (12c) or (12d), the probability of collisions in their respective dimensions increases substantially.
TABLE III: Process statistics
cap. Error(cm) cap. Time(s) cal. Time(s)
$0.98\pm 0.44$ $0.73\pm 0.12$ $0.10\pm 0.03$
TABLE IV: Collision avoidance statistics
Min Groud Dis.(m). Min Base Dis.(m)
Compliance $0.46\pm 0.20$ $0.29\pm 0.12$
TABLE V: Similarity statistics of trajectories
Intra-group Dis.(m) Inter-group Dis.(m) P Value
DTW Distance $8.39\pm 4.35$ $9.12\pm 4.72$ 0.61
TABLE VI: Comparison of peak torque
Primeval Compliance P Value
Max Peak Torque $0.87\pm 0.11$ $0.62\pm 0.08$ $091.17e-$ 09
V-D Real-World Scenarios
@@
As shown in Fig. 8, a real-world experiment is conducted on a mobile manipulator. We use a C-100 differential mobile base ($n_{b}$ = 2) and a franka panda manipulator ($n_{m}$ = 7), interfacing with the robot through ROS Noetic. NOKOV motion capture provides position and velocity measurements, which are used for both ball state estimation and mobile base state tracking. We use a standard tennis ball with a reflective strip (diameter: 6.8 cm, mass: 60.3 g) for tracking purposes, along with an inelastic mesh bag as the container which would not weaken the impact. Due to the coupling effects of differential wheel steering, our high-level planner considers only the combined motion $q_{1}$ and restricts the differential motion $q_{0}$ to 0 to achieve more accurate tracking in the PRC. When a human hand throws the ball, a Kalman filter models the ball’s dynamics using the initial frames and updates the model in the subsequent few frames. This model is then used to predict the ball’s trajectory over the next 1.5s. After that, the High-Level planner, PRE planner, and network inference will be executed sequentially, and each will only be executed once. It is worth noting that we tested the delay in ROS master-slave communication and the robot’s velocity control startup time. This delay was accounted for by sending the POC planning command in advance, allowing the robot to start the POC planning right around the time of the catch.
Figure 9: The change of the joint angles of the robot during a catch, the expected joint angles calculated by the high-level planner, and the trajectory of the ball and the container.
Figure 10: The curves of normalized joint torques. The maximum peak torque of all joints is 0.606.
We conducted a total of 100 hand-throwing experiments for compliant catching. We assess the effectiveness of the hand-thrown ball based on the closest distance between the container’s initial position and the ball’s trajectory. Due to inaccuracies in hand throwing, 19 attempts were deemed invalid. Out of the remaining 81 valid throws, the robot successfully captured the ball 75 times, resulting in a success rate of 92.59%, without any collision. The failure occurred because noise in the capture system caused a discrepancy between the filter’s estimated state and the actual state, leading to deviations in the predicted trajectory. It is worth mentioning that, due to the limitation and some obstacles of the laboratory room (some areas wound do not capture the ball well), we did not test in all directions like the simulation, but we still tried to throw to different directions of the robot (left, right, top and middle). In all experiments, TABLE III presents the error between the container’s capture point and the actual ball position ”cap. Error”, along with the capture time ”cap. Time” and the total calculation time ”cal. Time” for the Hive-Level, the PRC, and the Network inference phase. TABLE IV provides collision avoidance data, including the minimum distances between the container and the ground “Min Ground Dis.”, as well as the container and the mobile base ”Min Base Dis.”, during the POC phase. We successfully achieved a high precision capture and collision avoidance.
We compare the effect of adding compliance to reduce torque impact. Since different ball trajectories can result in different torque impacts, ensuring trajectory similarity in the comparison is crucial. To control for variables, we ask the person to consistently face the same direction to maintain similarity and record data 30 times for both the compliant and non-compliant conditions. In each of the 30 trajectories, we quickly select the 20 most similar ones by calculating the vector from the container’s initial position to the closest point on trajectory. We then use Dynamic Time Warping (DTW) to assess the similarity between the different group trajectories. The intra-group distance for the compliant group and the inter-group distance between the compliant and non-compliant groups have been calculated, as shown in TABLE V. The final p-value is much greater than the significance level of 0.05, indicating that the ball throwing trajectories of the two groups are similar. Therefore, the impact analysis can proceed to the next step.
Due to different hardware in each joint, every torque was normalized according to the maximum allowed torque at that joint, thus, ${\tau^{\prime}}_{i}=\tau_{i}/\tau_{i}^{\max},i\in[2,8]$. We calculate the maximum peak torque $max(\tau^{\prime}_{i})$ in all joints during the experiment, as the ”Max Peak Torque” in TABLE VI. The p-value for the ”Max Peak Torque” between the two groups is significantly lower than the 0.05 significance level, indicating a notable difference between the groups and confirming the feasibility of our compliant strategy. Using the mean value calculation, the quantitative attenuation ratio is $(0.87-0.62)/0.87=28.7\%$. Since there is currently no research specifically on compliant catching for mobile manipulators, we referred to the compliant ratio data for a fixed manipulator catching a vertically falling object [7], which is 28%. This result is similar to ours. However, our study is conducted in a more complex catching scenario and with a mobile manipulator. For one test in compliant catching, Figure 9(a) shows the expected output of the high-level capture planner, along with the changes in joint angles during the catching process. At the same time, Figure 9(b) presents the actual predicted trajectory of the ball and the container’s trajectory during the compliant catching experiment. Additionally, Fig. 10 illustrates the normalized joint torques.